OpenStudy (anonymous):

a sweater originally priced at $32 is on sale for $28.what is the percent of decrease rounded to the nearest whole number?

OpenStudy (anonymous):

Original price = $32 Sale price = $28 Decrease = $4 Decrease%=(Original price - Sale price)*100/(original price) Decrease%=(32-28)*100/32 Decrease%=4*100/32 Decrease%=400/32 Decrease%=12.5% Check ===== 12.5% of 32 = 32*(12.5/100)=400/100=4 Sale price = original price - discount Sale price = 32-4 = 28

OpenStudy (anonymous):

id round it to 13 :)